OpenStudy (anonymous):

6.2(-3x) + (-4.1)(-3x)

OpenStudy (anonymous):

is t 6.2 or 6*2?

OpenStudy (anonymous):

\[\frac{62}{10}*(-3x) - \frac{41}{10}*(-3x)=-\frac{63 x}{10} \]